# Huge Ensembles Part II: Properties of a Huge Ensemble of Hindcasts Generated with Spherical Fourier Neural Operators

### v2

@@ F.5 Aggregation
Appendix G Uncertainties in EFI, ROC, and twCRPS metrics due to finite sample size
The question of how uncertainties in the Extreme Forecast Index (EFI), Receiver Operating Characteristic (ROC), and twCRPS metrics scale with ensemble size can be answered by determining how the uncertainties in the empirical estimates of cumulative distribution functions (ECDFs), which govern all three metrics, scale with sample size relative to the "true" CDFs in the limit . The Dvoretzky–Kiefer–Wolfowitz (DKW) (Massart, 1990) inequality provides the mechanism for quantitative estimates of these uncertainties.
@@ Appendix G Uncertainties in EFI, ROC, and twCRPS metrics due to finite sample size
Actually, this limit is attained to within 1% for modest values of , specifically when .
@@ Acknowledgements.
- Bauer et al. (2015) Bauer, P., Thorpe, A., and Brunet, G.: The quiet revolution of numerical weather prediction, Nature, 525, 47–55, 10.1038/nature14956, 2015.
@@
- Gneiting and Ranjan (2011) Gneiting, T. and Ranjan, R.: Comparing Density Forecasts Using Threshold- and Quantile-Weighted Scoring Rules, Journal of Business & Economic Statistics, 29, 411–422, https://EconPapers.repec.org/RePEc:bes:jnlbes:v:29:i:3:y:2011:p:411-422, 2011.

### v3

@@ G.5 Aggregation
Appendix H Uncertainties in EFI, ROC, and twCRPS metrics due to finite sample size
The question of how uncertainties in the Extreme Forecast Index (EFI) and CRPS metrics scale with ensemble size can be answered by determining how the uncertainties in the empirical estimates of cumulative distribution functions (ECDFs), which govern all three metrics, scale with sample size relative to the "true" CDFs in the limit . The Dvoretzky–Kiefer–Wolfowitz (DKW) (Massart, 1990) inequality provides the mechanism for quantitative estimates of these uncertainties.
@@ Acknowledgements.
- Finkel et al. (2023) Finkel, J., Gerber, E. P., Abbot, D. S., and Weare, J.: Revealing the Statistics of Extreme Events Hidden in Short Weather Forecast Data, AGU Advances, 4, 10.1029/2023av000881, 2023.
@@
- Mahesh et al. (2024) Mahesh, A., Collins, W., Bonev, B., Brenowitz, N., Cohen, Y., Elms, J., Harrington, P., Kashinath, K., Kurth, T., North, J., OBrien, T., Pritchard, M., Pruitt, D., Risser, M., Subramanian, S., and Willard, J.: Huge Ensembles Part I: Design of Ensemble Weather Forecasts using Spherical Fourier Neural Operators, 10.48550/ARXIV.2408.03100, 2024.
@@
- Millin and Furtado (2022) Millin, O. T. and Furtado, J. C.: The Role of Wave Breaking in the Development and Subseasonal Forecasts of the February 2021 Great Plains Cold Air Outbreak, Geophysical Research Letters, 49, 10.1029/2022gl100835, 2022.
@@
- Steadman (1979) Steadman, R. G.: The Assessment of Sultriness. Part I: A Temperature-Humidity Index Based on Human Physiology and Clothing Science, Journal of Applied Meteorology, 18, 861–873, 10.1175/1520-0450(1979)018<0861:taospi>2.0.co;2, 1979.
@@
- Zamo and Naveau (2017) Zamo, M. and Naveau, P.: Estimation of the Continuous Ranked Probability Score with Limited Information and Applications to Ensemble Weather Forecasts, Mathematical Geosciences, 50, 209–234, 10.1007/s11004-017-9709-7, 2017.